Could someone please show me how to get the altitude of a triangle? I cant seem to remember the formula to solve. Please help me.
Side A 13
Side B 13
Bottom 10

Answer by edjones(8007)

The altitude divides the isoscoles triangle into 2 equal right triangles.
let a=altitude b=10/2=5 c=13
a^2+b^2=c^2
5^2+a^2=13^2
a^2=169-25=144
altitude=12
Ed